# Monte-Carlo Integration Based Multiple-Scattering Channel Modeling for Ultraviolet Communications in Turbulent Atmosphere

### v1

@@ II-B Absorption Effects in Turbulent Channels
[…]
@@ II-C Fluctuation Effects in Turbulent Channels
Besides the scattering and absorption effects, the turbulence also introduces a fluctuation effect on the light irradiance [45]. The turbulent fluctuation effect can be characterized by the turbulent fading coefficient $\eta$ with $\eta\triangleq P_{r}/P_{t}$ and $\quad\eta\geq 0$, where $P_{t}$ and $P_{r}$ are the transmit power and receiving power of a traveling process, respectively.
For weak turbulent conditions, the turbulent fading coefficient is usually modeled as an LN distribution with the normalized probability density function (PDF) given by [45]
@@
where $\sigma_{I}^{2}(d)=e^{\sigma_{r}^{2}(d)}-1$ and $\sigma_{r}^{2}(d)$ is the Rytov variance for a propagating distance $d$. For a plane wave, the Rytov variance is given by [45]
@@
For moderate and strong turbulent conditions, the turbulent fading coefficient is usually modeled as a GG distribution with the normalized PDF given by [45, 57]
$f_{H,GG}(\eta;d)$ $=\frac{2[\alpha(d)\beta(d)]^{\frac{\alpha(d)+\beta(d)}{2}}\eta^{\frac{\alpha(d)+\beta(d)}{2}-1}}{\Gamma(\alpha(d))\Gamma(\beta(d))}K_{\alpha(d)-\beta(d)}\left(2\sqrt{\alpha(d)\beta(d)\eta}\right),\quad\eta\geq 0,$ (13)
where $K_{v}(x)$ is the modified Bessel function of the second kind with parameter $v$; $\alpha(d)$ and $\beta(d)$ are related to the Rytov variance $\sigma_{r}^{2}(d)$ as [45, 57]
@@ III Receiving Power of Multiple-Scattering in Turbulent Channels
We first introduce the geometry setting for the multiple-scattering process of NLOS UV communications, shown in Fig. 1. The receiver (Rx) locates at the origin $(0,0,0)$ and the transmitter (Tx) locates at the $y$-axis with coordinates $(0,r,0)$. The direction cosines of Tx and Rx pointing directions are denoted by $\bm{\mu}_{T}$ and $\bm{\mu}_{R}$, respectively; The zenith angle and the azimuth angle of the Tx(Rx) are denoted by $\theta_{T(R)}$ and $\phi_{T(R)}$, respectively. The divergence angle of the Tx light beam is denoted by $\beta_{T}$ and the divergence angle of the Rx field-of-view (FOV) is denoted by $\beta_{R}$. We regard the initial photon emitting as the $zero$-order scattering and denote the propagating distance, scattering zenith angle and scattering azimuth angle of $i$-order scattering by $d_{i}$, $\theta_{i}$ and $\phi_{i}$, respectively. The position of the $i$th scatter is denoted by $\bm{r}_{i}$. The direction cosine of $i$-order scattering is denoted by $\bm{\mu}_{i}$. The initial scattering zenith angle $\theta_{0}$ is the angle between $\bm{\mu}_{0}$ and $\bm{\mu}_{T}$. We consider an $n$-order scattering process and define $\bm{d}\triangleq[d_{0},d_{1},\cdots,d_{n-1}]^{\text{T}}$, $\bm{\theta}\triangleq[\theta_{0},\theta_{1},\cdots,\theta_{n-1}]^{\text{T}}$, and $\bm{\phi}\triangleq[\phi_{0},\phi_{1},\cdots,\phi_{n-1}]^{\text{T}}$.
@@ III-A Receiving Power Ignoring Turbulent Fluctuation
A photon propagating path of an $n$-order scattering is full characterized by the parameters $\{\bm{d},\bm{\theta},\bm{\phi}\}$. After scattered by the last scatter, the photon can be detected if and only if the position of the $n$th scatter locates in the FOV of the receiver. We denote the set of parameters satisfying this geometrical constrain by $\mathcal{F}\triangleq\{\bm{d},\bm{\theta},\bm{\phi}|\bm{r}_{n}\text{ in FOV}\}=\{\bm{d},\bm{\theta},\bm{\phi}|\bm{r}_{n}\cdot\bm{\mu}_{R}\geq d_{n}\cos(\beta_{R}/2)\}$. Then the conditional receiving probability of an $n$-order scattering process for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be approximated as [17]
$P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ $=I_{n}(\bm{d},\bm{\theta},\bm{\phi})\left(\frac{k_{s}^{tot}}{k_{e}^{tot}}\right)^{n}\exp({-k_{e}^{tot}d_{n}})\cos\phi_{r}\min\left(1,p^{tot}(\theta_{r,n})\Omega_{r}\right),$ (15)
where $d_{n}\triangleq\|\bm{r}_{n}\|$ and $\theta_{r,n}$ is the angle between $\bm{\mu}_{n-1}$ and $-\bm{r}_{n}$; $\phi_{r}$ is the angle between $\bm{\mu}_{R}$ and $\bm{r}_{n}$; $\Omega_{r}$ is the solid angle formed by $\bm{r}_{n}$ and the receiving area $A_{r}$; and we have $\Omega_{r}\approx 2\pi(1-d_{n}/\sqrt{d_{n}^{2}+r_{A}^{2}})$, where $r_{A}=\sqrt{A_{r}/\pi}$ is the radius of the receiving area; $I_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is an indicator function; and we have $I_{n}(\bm{d},\bm{\theta},\bm{\phi})=1$ and $I_{n}(\bm{d},\bm{\theta},\bm{\phi})=0$ when the $n$th scatter locates within and without the FOV, respectively. Then the conditional receiving probability can be rewritten as
@@
$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (17)
where $f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$. We have used the fact that $\bm{d}$, $\bm{\theta}$, and $\bm{\phi}$ are independent from each other. For the propagating distance $d_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{D}(d_{i})=k_{e}^{tot}e^{-k_{e}^{tot}d_{i}}$ with $0\leq d_{i}\leq\infty$ [17]. For the scattering angle $\theta_{i}$ with $i=0$, we have $f_{\Theta}(\theta_{0})=\frac{\sin\theta_{0}}{1-\cos({\beta_{T}}/{2})}$ with $0\leq\theta_{0}\leq{\beta_{T}}/{2}$ when a uniform light source is assumed [17]; and for the scattering angle $\theta_{i}$ with $i\in\{1,2,\cdots,n-1\}$, we have $f_{\Theta}(\theta_{i})=2\pi p^{tot}(\theta_{i})\sin\theta_{i}$ with $0\leq\theta_{i}\leq\pi$ [17]. For the scattering azimuth angle $\phi_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{\Phi}(\phi_{i})=\frac{1}{2\pi}$ with $0\leq\phi_{i}\leq 2\pi$ [17].
@@ III-B Receiving Power Considering Turbulent Fluctuation
The conditional receiving probability without turbulent fluctuation for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ is $P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ given in (15). Then the conditional receiving probability under turbulent fluctuation effect for a given set of turbulent fading coefficients $\bm{\eta}\triangleq[\eta_{0},\eta_{1},\cdots,\eta_{n}]^{\text{T}}$ and a propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be expressed as
$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n}\eta_{i}P_{n}(\bm{d},\bm{\theta},\bm{\phi}).$ (18)
@@
$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=\int_{\bm{\eta}}P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{\eta},$ (19)
where $f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$ conditioned on a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$. For a given propagating path, $\eta_{0},\eta_{1},\cdots,\eta_{n}$ are assumed as independent variables. Then we have
$f_{H|D,\Theta,\Phi}(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n}f_{H}(\eta_{i};d_{i}),$ (20)
where $f_{H}(\eta;d)=f_{H,LN}(\eta;d)$ is given in (11) for LN distribution and $f_{H}(\eta;d)=f_{H,GG}(\eta;d)$ is given in (13) for GG distribution.
@@
$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=P_{n}(\bm{d},\bm{\theta},\bm{\phi})\prod_{i=0}^{n}\left[\int_{\eta_{i}}\eta_{i}f_{H}(\eta_{i};d_{i})\mathrm{d}\eta_{i}\right]=P_{n}(\bm{d},\bm{\theta},\bm{\phi}),$ (21)
where we have used the property that $\int_{\eta_{i}}\eta_{i}f_{H}(\eta_{i};d_{i})\mathrm{d}\eta_{i}=1$ for $i=0,1,\cdots,n$.
@@
$P_{tur,n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=P_{n}.$ (22)
@@
IV Estimating Turbulent Channels Using MCI Approach
@@ IV-A Estimating the Average Receiving Power
The integration in (17) can be estimated by using the Monte-Carlo methods. In this paper, we use the MCI approach to estimate the average receiving power in (17). Compared with another widely used Monte-Carlo simulation (MCS) approach [13, 14], the MCI approach [15, 16, 17, 18, 19] enjoys a more flexible sampling function, resulting in a higher computational efficiency. Specifically, we can rewrite (17) as
@@
where $g_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is defined as $g_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq P_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})$, and where $P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is given in (16). Now we can choose a sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ and further rewrite the receiving power $P_{n}$ in (23) as
$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}O_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{n}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=\text{E}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})],$ (24)
where $O_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq g_{n}(\bm{d},\bm{\theta},\bm{\phi})/f_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ is called the objective function of MCI approach; and $\text{E}[x]$ denotes the expectation of $x$. Therefore, we have expressed the receiving power $P_{n}$ as the expectation of the objective function $O_{n}(\bm{d},\bm{\theta},\bm{\phi})$ when the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$ is the sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$.
Different sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ can result in different computational efficiency of MCI approach [15, 16, 17]. Here we adopt the so-called important sampling method [15], which enjoys the same convergent speed as the MCS approach. The sampling PDF of the important sampling is given by
$f_{n}(\bm{d},\bm{\theta},\bm{\phi})=f_{D,\Theta,\Phi}(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i}).$ (25)
@@
where $\text{rand}(1)$ is a random number between 0 and 1; $F_{\Theta}^{-1}(\theta_{s})$ is the inverse function of the cumulative distribution function of the scattering angle, i.e., $F_{\Theta}(\theta_{s})=\int_{0}^{\theta_{s}}f_{\Theta}(\theta)\mathrm{d}\theta$, which can be obtained as
$F_{\Theta}(\theta_{s})$ $=C_{1}\cos\theta_{s}+C_{2}\cos^{3}\theta_{s}+C_{3}+C_{4}(C_{5}-C_{6}\cos\theta_{s})^{-1/2}+\frac{C_{7}(1-\cos\theta_{s})}{2+C_{8}(1-\cos\theta_{s})},$ (29)
where $C_{1}=-\frac{3(1+3\gamma)}{8(1+2\gamma)}\frac{k_{s}^{ray}}{k_{s}^{tot}}+\frac{f(1-g^{2})}{4(1+g^{2})^{3/2}}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{2}=-\frac{1-\gamma}{8(1+2\gamma)}\frac{k_{s}^{ray}}{k_{s}^{tot}}-\frac{f(1-g^{2})}{4(1+g^{2})^{3/2}}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{3}=\frac{1}{2}\frac{k_{s}^{ray}}{k_{s}^{tot}}+\frac{1+g}{2g}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{4}=-\frac{1-g^{2}}{2g}\frac{k_{s}^{mie}}{k_{s}^{tot}}$, $C_{5}=1+g^{2}$, $C_{6}=2g$, $C_{7}=(1+4k^{2}d_{0}^{2})\frac{k_{s}^{tur}}{k_{s}^{tot}}$, and $C_{8}=4k^{2}d_{0}^{2}$. For $i=1,2,\cdots,n-1$, the scattering angle $\theta_{i}$ can be solved by using numerical methods, e.g., Newton’s bisection search.
@@
$O_{n}(\bm{d},\bm{\theta},\bm{\phi})$ $=P_{n}(\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}P_{d,n},\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\in\mathcal{F},\\ 0,\quad\{\bm{d},\bm{\theta},\bm{\phi}\}\notin\mathcal{F}.\end{cases}$ (30)
@@
Then we can approximate the receiving power when $M$ is large as
@@
IV-B Estimating the Channel Impulse Response
For each sampling point $\bm{s}_{n}^{m}$, we can calculating the corresponding receiving time as
$t_{n}^{m}=\frac{\sum_{i=0}^{n}d_{i}^{m}}{c},$ (34)
where $c$ is the light speed in the atmosphere. Then according to the result in [17], we can obtain the CIR of the multiple-scattering of UV communications in the following steps:
  • Step 1: Classify the objective function value $O_{n}(\bm{s}_{n}^{m})$ into $L$ different time slots $\{(0,\Delta t)$, $(\Delta t,2\Delta t),\cdots,((L-1)\Delta t,L\Delta t)\}$ according to its receiving time $t_{n}^{m}$, where $\Delta t$ is a pre-chosen time interval;
  • Step 2: Add up objective function values in each time slot and divide it by the number of sampling points $M$. Then we obtain the receiving probability sequences $[{P}_{n}(\Delta t),{P}_{n}(2\Delta t)$, $\cdots,{P}_{n}(L\Delta t)]^{\text{T}}$;
  • Step 3: The CIR is obtained by normalizing the sequences $[{P}_{n}(\Delta t),{P}_{n}(2\Delta t),\!\cdots,\!{P}_{n}(L\Delta t)]^{\text{T}}$ using the receiving time interval $\Delta t$ and the receiving area $A_{r}$. Therefore, the value of CIR in a unit of irradiance at $t=l\Delta t$ is given by
    $h_{n}(l\Delta t)=\frac{P_{n}(l\Delta t)}{\Delta_{t}A_{r}},\quad l=1,2\cdots,L.$ (35)
Then the total CIR at $t=l\Delta t$ over $N$ scattering orders can be obtained as
$h_{tot}(l\Delta t)=\sum_{n=1}^{N}\frac{P_{n}(l\Delta t)}{\Delta_{T}A_{r}},\quad l=1,2\cdots,L.$ (36)
IV-C Estimating the Turbulent Fluctuation
It is challenging to estimate the turbulent fluctuation of the receiving power by using Monte-Carlo based methods. This is because the Monte-Carlo process itself will introduce a fluctuation effect on the calculated average receiving power [17]. In existing literature [52], the turbulent fluctuation was estimated by simply calculating the variance of the average receiving power $P_{tur,n}$, i.e, $\text{Var}[P_{tur,n}]$, obtained from the Monte-Carlo models. However, because $P_{tur,n}=P_{n}$, using the expression in (32), we can obtain

### v3

@@ II-B Absorption Effects in Turbulent Channels
Besides the scattering effects, the light wave can also be selectively absorbed by the particles due to the electronic transition between different energy levels. The absorption coefficient caused by the turbulence depends on the complex dielectric constant $\epsilon=\epsilon_{r}+i\epsilon_{i}$, i.e., $k_{a}^{tur}=k\epsilon_{i}\epsilon_{r}^{-\frac{1}{2}}$, where $\epsilon_{i}=60\lambda\delta$ and $\delta$ is the atmospheric conductivity. For the atmosphere near the ground, we have $\epsilon_{r}\approx 1.00059$ and $\delta\approx 2.2\times 10^{-14}$ $\text{S}/\text{m}$. Then we can obtain the absorption coefficient due to the turbulence as $k_{a}^{tur}\approx 4.1457\times 10^{-12}$, which is much smaller than the absorption coefficient due to the particles. Therefore, we can always ignore the absorption effect due to the turbulence for UV communications.
@@ II-C Fluctuation Effects in Turbulent Channels
Besides the scattering and absorption effects, the turbulence also introduces a fluctuation effect on the light irradiance [andrews2005laser]. The turbulent fluctuation effect can be characterized by the turbulent fading coefficient $\eta$ with $\eta\triangleq P_{r}/P_{t}$ and $\eta\geq 0$, where $P_{t}$ and $P_{r}$ are the transmit power and receiving power of a traveling process, respectively.
For weak turbulent conditions, the turbulent fading coefficient is usually modeled as an LN distribution with the normalized probability density function (PDF) given by [andrews2005laser]
@@
where $\sigma_{I}^{2}(d)=e^{\sigma_{r}^{2}(d)}-1$ and $\sigma_{r}^{2}(d)$ is the Rytov variance for a propagating distance $d$. For a plane wave, the Rytov variance is given by [andrews2005laser]
@@
For moderate and strong turbulent conditions, the turbulent fading coefficient is usually modeled as a GG distribution with the normalized PDF given by [andrews2005laser]
$f_{H,GG}(\eta;d)$ $=\frac{2[\alpha(d)\beta(d)]^{\frac{\alpha(d)+\beta(d)}{2}}\eta^{\frac{\alpha(d)+\beta(d)}{2}-1}}{\Gamma(\alpha(d))\Gamma(\beta(d))}$ (10)
$\quad\quad\times K_{\alpha(d)-\beta(d)}\left(2\sqrt{\alpha(d)\beta(d)\eta}\right),\quad\eta\geq 0,$
where $K_{v}(x)$ is the modified Bessel function of the second kind with parameter $v$; $\alpha(d)$ and $\beta(d)$ are related to the Rytov variance $\sigma_{r}^{2}(d)$ as [andrews2005laser]
@@ III Receiving Power of Multiple-Scattering in Turbulent Channels
The geometry setting for the multiple-scattering process of NLOS UV communications is shown in Fig. 1. The receiver (Rx) locates at the origin $(0,0,0)$ and the transmitter (Tx) locates at the $y$-axis with coordinates $(0,r,0)$. The direction cosines of Tx and Rx pointing directions are denoted by $\bm{\mu}_{T}$ and $\bm{\mu}_{R}$, respectively. The zenith angle and the azimuth angle of the Tx (Rx) are denoted by $\theta_{T(R)}$ and $\phi_{T(R)}$, respectively. The divergence angle of the Tx light beam is denoted by $\beta_{T}$ and the divergence angle of the Rx field-of-view (FOV) is denoted by $\beta_{R}$. We regard the initial photon emitting as the zerozero-order scattering and denote the propagating distance, scattering zenith angle and scattering azimuth angle of $i$-order scattering by $d_{i}$, $\theta_{i}$ and $\phi_{i}$, respectively. The position of the $i$th scatter is denoted by $\bm{r}_{i}$. The direction cosine of $i$-order scattering is denoted by $\bm{\mu}_{i}$. The initial scattering zenith angle $\theta_{0}$ is the angle between $\bm{\mu}_{0}$ and $\bm{\mu}_{T}$. We consider an $n$-order scattering process and define $\bm{d}\triangleq[d_{0},d_{1},\cdots,d_{n-1}]^{\text{T}}$, $\bm{\theta}\triangleq[\theta_{0},\theta_{1},\cdots,\theta_{n-1}]^{\text{T}}$, and $\bm{\phi}\triangleq[\phi_{0},\phi_{1},\cdots,\phi_{n-1}]^{\text{T}}$.
@@ III-A Receiving Power Ignoring Turbulent Fluctuation
A photon propagating path of an $n$-order scattering is full characterized by the parameters $\{\bm{d},\bm{\theta},\bm{\phi}\}$. After scattered by the last scatter, the photon can be detected if and only if the position of the $n$th scatter locates in the FOV of the receiver. We denote the set of parameters satisfying this geometrical constrain by $\mathcal{F}\triangleq\{\bm{d},\bm{\theta},\bm{\phi}|\bm{r}_{n}\text{ in FOV}\}=\{\bm{d},\bm{\theta},\bm{\phi}|\bm{r}_{n}\cdot\bm{\mu}_{R}\geq d_{n}\cos(\beta_{R}/2)\}$. Then the conditional receiving probability of an $n$-order scattering process for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be approximated as [yuan2019monte]
$P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ $=I_{n}(\bm{d},\bm{\theta},\bm{\phi})\left(\frac{k_{s}^{tot}}{k_{e}^{tot}}\right)^{n}\exp({-k_{e}^{tot}d_{n}})$ (12)
$\quad\quad\times\cos\phi_{r}\min\left(1,p^{tot}(\theta_{r,n})\Omega_{r}\right),$
where $d_{n}\triangleq\|\bm{r}_{n}\|$ and $\theta_{r,n}$ is the angle between $\bm{\mu}_{n-1}$ and $\bm{r}_{n}$; $\phi_{r}$ is the angle between $\bm{\mu}_{R}$ and $\bm{r}_{n}$, as shown in Fig. 1; $\Omega_{r}$ is the solid angle formed by $\bm{r}_{n}$ and the receiving area $A_{r}$; and we have $\Omega_{r}\approx 2\pi(1-d_{n}/\sqrt{d_{n}^{2}+r_{A}^{2}})$, where $r_{A}=\sqrt{A_{r}/\pi}$ is the radius of the receiving area; and $I_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is an indicator function. The indicator function $I_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is defined as
$I_{n}(\bm{d},\bm{\theta},\bm{\phi})=\begin{cases}1,&\text{$\bm{r_{n}}$ locates in FOV},\\ 0,&\text{otherwise}.\end{cases}$ (13)
Then the conditional receiving probability can be rewritten as
@@
$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi},$ (15)
where $f(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i})$ is the joint PDF of $\{\bm{d},\bm{\theta},\bm{\phi}\}$. We have used the fact that $\bm{d}$, $\bm{\theta}$, and $\bm{\phi}$ are independent from each other. For the propagating distance $d_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{D}(d_{i})=k_{e}^{tot}e^{-k_{e}^{tot}d_{i}}$ with $0\leq d_{i}\leq\infty$ [yuan2019monte]. For the scattering angle $\theta_{i}$ with $i=0$, we have $f_{\Theta}(\theta_{0})=\frac{\sin\theta_{0}}{1-\cos({\beta_{T}}/{2})}$ with $0\leq\theta_{0}\leq{\beta_{T}}/{2}$ when a uniform light source is assumed [yuan2019monte]; and for the scattering angle $\theta_{i}$ with $i\in\{1,2,\cdots,n-1\}$, we have $f_{\Theta}(\theta_{i})=2\pi p^{tot}(\theta_{i})\sin\theta_{i}$ with $0\leq\theta_{i}\leq\pi$ [yuan2019monte]. For the scattering azimuth angle $\phi_{i}$ with $i\in\{0,1,\cdots,n-1\}$, we have $f_{\Phi}(\phi_{i})=\frac{1}{2\pi}$ with $0\leq\phi_{i}\leq 2\pi$ [yuan2019monte].
@@ III-B Receiving Power Considering Turbulent Fluctuation
The conditional receiving probability without turbulent fluctuation for a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ is $P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ given in (12). Then the conditional receiving probability under turbulent fluctuation effect for a given set of turbulent fading coefficients $\bm{\eta}\triangleq[\eta_{0},\eta_{1},\cdots,\eta_{n}]^{\text{T}}$ under a propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$ can be expressed as
$P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})=P_{n}(\bm{d},\bm{\theta},\bm{\phi})\prod_{i=0}^{n}\eta_{i}.$ (16)
@@
$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=\int_{\bm{\eta}}P_{tur,n}(\bm{\eta},\bm{d},\bm{\theta},\bm{\phi})f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{\eta},$ (17)
where $f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})$ is the joint PDF of $\{\eta_{0},\eta_{1},\cdots,\eta_{n}\}$ conditioned on a given propagating path $\{\bm{d},\bm{\theta},\bm{\phi}\}$. For a given propagating path, $\eta_{0},\eta_{1},\cdots,\eta_{n}$ are assumed as independent variables. Then we have
$f(\bm{\eta}|\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n}f(\eta_{i};\sigma_{I}^{2}(d_{i})),$ (18)
where $f(\eta;d)=f_{H,LN}(\eta;d)$ is given in (8) for LN distribution and $f(\eta;d)=f_{H,GG}(\eta;d)$ is given in (10) for GG distribution.
@@
$P_{tur,n}(\bm{d},\bm{\theta},\bm{\phi})$ $=P_{n}(\bm{d},\bm{\theta},\bm{\phi})\prod_{i=0}^{n}\left[\int_{\eta_{i}}\eta_{i}f(\eta_{i};\sigma_{I}^{2}(d_{i}))\mathrm{d}\eta_{i}\right]$ (19)
$=P_{n}(\bm{d},\bm{\theta},\bm{\phi}),$
where we have used the property $\int_{\eta_{i}}\eta_{i}f(\eta_{i};\sigma_{I}^{2}(d_{i}))\mathrm{d}\eta_{i}=1$ for $i=0,1,\cdots,n$.
@@
$P_{tur,n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}P_{n}(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}=P_{n}.$ (20)
@@
IV Estimating Turbulent Channels by Using an MCI Approach
@@ IV-A Estimating the Average Receiving Power
The integration in (15) can be estimated by using an MCI approach. Compared with another widely used MCS approach [ding2009modeling, drost2011uv], the MCI approach [yuan2019importance, ding2010path, yuan2019monte, shan2020modeling, shen2020improved] enjoys a more flexible sampling function, resulting in a higher computational efficiency. Besides, as we will demonstrate later in Section IV-B, the MCI approach can provide a more interpretable logic for estimating the turbulent fluctuation.
Following the MCI procedure given in [yuan2019monte], we can rewrite (15) as
@@
where $g_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is defined as $g_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq P_{n}(\bm{d},\bm{\theta},\bm{\phi})f(\bm{d},\bm{\theta},\bm{\phi})$, and where $P_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is given in (14). Now we can choose a sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ and further rewrite the receiving power $P_{n}$ in (21) as
$P_{n}$ $=\int_{\bm{d}}\int_{\bm{\theta}}\int_{\bm{\phi}}O_{n}(\bm{d},\bm{\theta},\bm{\phi})f_{n}(\bm{d},\bm{\theta},\bm{\phi})\mathrm{d}\bm{d}\mathrm{d}\bm{\theta}\mathrm{d}\bm{\phi}$ (22)
$=\text{E}[O_{n}(\bm{d},\bm{\theta},\bm{\phi})],$
where $O_{n}(\bm{d},\bm{\theta},\bm{\phi})\triangleq g_{n}(\bm{d},\bm{\theta},\bm{\phi})/f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is called the objective function of MCI approach; and $\text{E}[x]$ denotes the expectation of $x$. Therefore, we have expressed the receiving power $P_{n}$ as the expectation of the objective function $O_{n}(\bm{d},\bm{\theta},\bm{\phi})$ when a sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ is used to sample $\{\bm{d},\bm{\theta},\bm{\phi}\}$.
Different sampling PDF $f_{n}(\bm{d},\bm{\theta},\bm{\phi})$ can result in different computational efficiency of MCI approach [yuan2019importance, ding2010path, yuan2019monte]. Here we adopt the so-called important sampling method [yuan2019importance], which enjoys the same convergent speed as the MCS approach but a faster calculating speed. The sampling PDF of the important sampling is given by
$f_{n}(\bm{d},\bm{\theta},\bm{\phi})=f(\bm{d},\bm{\theta},\bm{\phi})=\prod_{i=0}^{n-1}f_{D}(d_{i})f_{\Theta}(\theta_{i})f_{\Phi}(\phi_{i}).$ (23)
@@
where $\text{rand}(1)$ is a random number between 0 and 1; $F_{\Theta}^{-1}(\theta_{s})$ denotes the inverse cumulative distribution function of the scattering angle $\theta_{s}$. For $i=1,2,\cdots,n-1$, the scattering angle $\theta_{i}$ can be solved by using numerical methods, e.g., Newton’s bisection search.
@@
[…]
@@
Then when $M$ is large, we can approximate the receiving power as
@@
IV-B Estimating the Turbulent Fluctuation
It is challenging to estimate the turbulent fluctuation of the receiving power by using Monte-Carlo based methods. This is because the Monte-Carlo process itself will introduce a fluctuation effect on the calculated average receiving power [yuan2019monte]. In existing literature [wang2013characteristics], the turbulent fluctuation was estimated by simply calculating the variance of the average receiving power $P_{tur,n}$, i.e, $\text{Var}[P_{tur,n}]$, obtained from the Monte-Carlo models. However, because $P_{tur,n}=P_{n}$, using the expression in (29), we can obtain